# Denominators in Lusztig’s asymptotic Hecke algebra via the Plancherel formula

### v5

@@ 2.5 The Plancherel formula for $\mathrm{GL}_{n}$
$f(i_{M})=\sigma(i_{M})\delta^{\frac{1}{2}}_{P}(i_{M})f(1)=f(1),$
and so if $\delta^{\frac{1}{2}}_{P}(i_{M})=1$, then it must be that $f(1)\in\sigma^{I_{M}}$. The Iwahori factorization
$I_{M}=N_{P}^{-}(\pi\mathcal{O})A(\mathcal{O})N_{M}(\mathcal{O})$
and the fact that $\mathrm{Ad}_{\mathfrak{n}_{P}}$ is a representation implies that
$\delta^{\frac{1}{2}}_{P}(i_{M})=|\det\mathrm{Ad}_{\mathfrak{n}_{P}}(N_{P}^{-}(i_{M})A(i_{M})N_{M}(i_{M}))|_{F}=|\det\mathrm{Ad}_{\mathfrak{n}_{P}}(A(i_{M}))|_{F}=1$
by unipotence of the groups $N_{P}^{-}$ and $N_{M}$ and the fact that the image of $i_{M}$ in $M(\mathbb{F}_{q})$ lies in the Borel. ∎
@@
Let $\pi:=\operatorname{Ind}_{P}^{G}(\nu\otimes\omega)$ be a tempered irreducible representation of $G$ induced as usual from a standard parabolic subgroup $P$ with Levi subgroup $M$. Then $\pi^{I}$ is a tempered irreducible $H$-module, and is of the form $K(u,s,\rho,q)$ for $u,s\in\mathbf{G}^{\vee}(\mathbb{C})$ and $\rho$ a representation of $\pi_{0}(Z_{G^{\vee}}(u,s))$ with $s$ compact. By [BK], Corollary 2.6, $K(u,s,\rho,q)$ extends to a $J$-module. By definition of Lusztig’s bijection, $\pi(f_{w})\neq 0$ only if $w$ is in the two-sided cell $\mathbf{c}=\mathbf{c}(u)$ of $\tilde{W}$ corresponding to $u$. On the other hand, $K(u,s,\rho,q)$ is induced from a square-integrable standard module $K_{M}(u,s,\tilde{\rho},q)$ of $H(M,I_{M})$. But now [KLDeligneLanglands, Theorem 8.3] says that $(u,s)$ must be preicsely as in the statement of the proposition. Thus only such summands $(f_{w})_{M}$ in the Plancherel decomposition of $f_{w}$ are nonzero. ∎
@@ 2.8.2 Conventions on parabolic subgroups
where we view $\beta_{i}$ as restricted to $\mathfrak{a}_{P}\mathrel{\hookrightarrow}\mathfrak{a}_{G}$. Note that all the positive roots $\Sigma^{+}(\theta)$ of $(P_{\theta},N_{\theta})$ are in $N_{P}$. Denoting by $[\alpha]$ the coset representing a root $\alpha$ of $G$ restricted to $P$, the positive roots in $N_{P}$ are the $\alpha_{ij}$ such that $[\alpha_{ij}]=[\beta_{n_{1}+\cdots+n_{k}}]$.
@@ 2.8.3 Regularity of the trace
That is, the trace is a regular function on (×$(\mathbb{C}^{\times})^{k}$.
@@
When $\nu$ is non-strictly positive with respect to $M$, by Theorem 7, $\mathrm{trace}\left({\pi}\,,{f}\right)$ has poles only of the form $z_{i}^{n_{i}}=0$. The claim now follows from the $W$-invariance, and thus $\mathrm{trace}\left({\pi}\,,{f}\right)$ is a regular function on (×$(\mathbb{C}^{\times})^{k}$. ∎
@@
Let $d\in\tilde{W}$ be a distinguished involution and $\pi$ be a tempered representation induced from one of the Levi subgroups attached to the two-sided cell containing $d$ in the sense of Proposition 1. Then $\mathrm{trace}\left({\pi}\,,{f_{d}}\right)$ is constant and a natural number. In fact, the same is true for any idempotent in $J$.
@@
3 Proof of Theorem 13 for $\mathrm{GL}_{n}$, and of Theorem 2 for general $\mathbf{G}$
@@ 3.1 The functions $f_{w}$ for $\mathrm{GL}_{n}$
To compute with the Plancherel formula, we will need to apply the residue theorem successively in each variables $z_{i}$, and in doing so will we will need to sum over a certain tree that will track, for each variable, at which residues we evaluated. Upon integrating with respect to each variable $z_{i}$, we will have poles of the form $z_{i}=0$ or $z_{i}=(q^{ij})^{-1}z_{j}$. For example, if we have $4$ variables $z_{1},z_{2},z_{3},z_{4}$ corresponding to a Levi subgroup $\mathrm{GL}_{l_{1}}\times\mathrm{GL}_{l_{2}}\times\mathrm{GL}_{l_{3}}\times\mathrm{GL}_{l_{4}}$, then some of the summands obtained by successively applying the residue theorem are labelled by paths on the tree
@@
Given a Levi subgroup $M$ with $N+1$ blocks, a bookkeeping tree $T$ for $M$ is a rooted tree with $N$ levels such that the vertices on the $i$-th level below the root each have $N-i$ child vertices, and each vertex is decorated with an equation of the form $z_{i}=0$ or $z_{i}=(q^{ij})^{-1}z_{j}$, where the index $j$ does not appear along the path from the vertex to the root, and the parent root is decorated with an equation $z_{k}=(q^{ki})^{-1}z_{i}$ for some $k$. Moreover, we require that the root be decorated with an equation of the form $z_{i}=0$ or $z_{i}=(q^{ij})^{-1}z_{j}$ for $i$ minimal. A branch of $T$ is a simple path in $T$ from the root to one of the leaves.
@@
We will show that each of the three forms of denominator that appear in the conclusion of Theorem 11 divide $P_{\mathbf{G}/\mathbf{B}}(q)$, and thus that their product divides a power of $P_{\mathbf{G}/\mathbf{B}}(q)$. The denominators of $c_{M}$ are all of the form $1+q+\cdots+q^{l_{i}-1}$, and so divide $P_{\mathbf{G}/\mathbf{B}}(q)$ as $l_{i}\leq n$ for all $i$. Note that as $l_{i_{i}}+l_{i_{k+1}}\leq n$, the leftmost denominators in (10) satisfy the conclusion of the corollary also. Finally, $Q_{rk}q^{i_{r}i_{k+1}}=q^{l_{i_{k+1}}+l_{i_{k}}+\cdots+l_{i_{r}}}$, and $R_{rk}$ is likewise always a polynomial in $q$ (as opposed to $q^{-1}$) divisible by $1-q$. Again using that $l_{i_{k+1}}+l_{i_{k}}+\cdots+l_{i_{r}}\leq n$, we are done with the first statement.
@@
for tempered principal series. Now let $w\in\mathbf{c}_{0}$. By Lemma 4 and the Satake isomorphism, there is a function $h_{w}$ in the spherical Hecke algebra such that for all principal series $\pi$,
@@
$f_{w}(1)=\int{\mathrm{trace}\left(\pi_{I}(f_{w})\right)}\,\mathrm{d}{\pi_{\mu_{I}}}=\int{\mathrm{trace}\left(\pi_{K}(h_{w})\right)}\,\mathrm{d}{\pi_{\mu_{I}}}=\frac{1}{P_{\mathbf{G}/\mathbf{B}}(q)}\int{\mathrm{trace}\left(\pi_{K}(h_{w})\right)}\,\mathrm{d}{\pi_{\mu_{K}}}=\frac{h_{w}(1)}{P_{\mathbf{G}/\mathbf{B}}(q)}.$ (12)
@@
Here the rank is given by [XiI], Proposition 5.5. (In op. cit. there is the assumption of simple-connectedness, but it is easy to see that the distinguished involutions for the extended affine Weyl group $\tilde{W}(\tilde{\mathbf{G}}(F))$ of the universal cover $\tilde{\mathbf{G}}$ are distinguished involutions for $\tilde{W}$ using the definition in [affineII] and uniqueness of the $\{C_{w}\}$-basis, and that the lowest cell is just the lowest cell of $\tilde{W}(\tilde{\mathbf{G}}(F))$ intersected with $\tilde{W}$.) ∎
@@
Let $w\in\tilde{W}$. Then $f_{w}(1)$ is a rational function of $q$ with denominator dividing a power of the Poincaré polynomial of $\mathbf{G}$. The numerator is a Laurent polynomial $p_{1}(q^{1/2})+p_{2}(q^{-1/2})$ in $q^{1/2}$, where the degree of $p_{1}$ is bounded uniformly in terms of $\tilde{W}$. The denominator of $f_{w}(1)$ depends only on the two-sided cell containing $w$.
@@
where we view $\nu$ as a semisimple conjugacy class in $M_{\mathbf{c}}$, and $V(\lambda)$ is the irreducible representation of $M_{\mathbf{c}}$ of highest weight $\lambda$. Then we have that $f_{\lambda}(1)\neq 0$ only if $\lambda$ is of height 00 with respect to the basis $\varepsilon_{i}\colon\mathrm{diag}(a_{1},\dots,a_{n})\mapsto a_{i}$.
@@ 3.1.2 Proofs of Theorem 11, Lemma 6, and Corollary 2
which is one of the factors in the product $(1-Q_{r,k-1}q_{i_{r}i_{k}})(Q_{r,k-1}(q_{i_{r}i_{k}})^{-1})$. The surviving factor in the numerator at index $(k-1,r)$ is then equal to $(1-Q_{r,k-1}q^{g_{i_{r}}-g_{i_{k}}})$. In short, the above factors in the denominator cancel with a numerator occurring with the same $r$-index but $k$-index one lower. Such a factor occurs whenever $r<k-1$ (note that this inequality does not hold when $k=1$ and $r=0$). When $r=k-1$, we have
@@
When evaluating (14) along a single branch, we find again that the only poles that appear are of the form $z_{i}=(q^{ij})^{-1}z_{j}$, or $z_{i}^{e_{i}}=0$ for $e_{i}<0$. We will track the effect that evaluating each successive residue has on the total degree of the integrand, and then conclude using the fact that $\int_{\mathbb{T}}z^{r}\,\mathrm{d}z=2\pi i\delta_{r,-1}$. First, observe that evaluating a residue of the form $z_{i}=(q^{ij})^{-1}z_{j}$ increases the sum of all powers by $1$; a factor $z_{j}$ is contributed to the resulting integrand. The sum of all powers is likewise increased by $1$ when evaluating the residue at a simple pole of $z_{i}^{-1}$ at 00. To compute the residue at a pole of $z_{i}^{-e_{i}}$ at 00 for $e_{i}>1$, consider the Taylor expansion at 00 of $\Gamma^{ij}$. We have
@@
$1+\frac{q^{ij}+(q^{ij})^{-1}-q_{ij}-q_{ij}^{-1}}{z_{j}}z_{i}+\frac{(q^{ij})^{2}+2+(q^{ij})^{-2}-(q_{ij}+q_{ij}^{-1})(q^{ij}+q^{-ij})}{z_{j}}z_{i}^{2}+\cdots\\ +\frac{-(q_{ij}+q_{ij}^{-1})((q^{ij})^{n-1}+\cdots+(q^{ij})^{-n+1})+(q^{ij})^{n-2}+\cdots+(q^{ij})^{-n+2}+(q^{ij})^{n}+\cdots+(q^{ij})^{-n}}{z_{j}^{n}}z_{i}^{n}\\ +\cdots.$ (17)
It is clear that the salient point of (17), that $z_{i}^{n}$ appears with a coefficient proportional to $z_{j}^{-n}$, holds also for the all the products like those in (16). Thus computing a residue at the pole of $z_{i}^{-e_{i}}$ at 00 will result in a new integrand, the total degree of which has increased by $e_{i}-e_{i}+1=1$. It now follows that after integrating with respect to $N-1$ variables, the final integral to be computed will be a constant times $(2\pi i)^{-1}\int_{\mathbb{T}}z_{N}^{e_{0}+\cdots+e_{N}+N-1}\mathrm{d}z_{N}$. This is nonzero if and only if $e_{0}+\cdots+e_{N}=-N$. Summing over all branches of a bookkeeping tree, we see that (14) is nonzero only if $e_{0}+\cdots+e_{N}=-N$. ∎
@@
The total degree of the integrand is now $e_{0}+\cdots+e_{N}+1=-(N-1)$. Therefore either $e_{i_{0}}+e_{i_{1}}+1=e_{i_{2}}=\cdots=e_{-N}=-1$, or we may again assume without loss of generality that the exponent of some $z_{i_{j}}$ is nonnegative, and proceed with evaluating (18) by integrating with respect to $z_{i_{j}}$. We may continue in this way, never having to deal with more than a simple pole at 00. (Of course, the resulting order of integration need not be the same as the order used in the proofs of Theorem 11 and Corollary 2.)
@@
for new exponents $e^{\prime}_{i_{k}}$. This and the calculations in the proof of Theorem 11 make clear that the only positive powers of $q^{1/2}$ that appear in the any denominator are those that appeared as in Theorem 11. These are controlled by the possible block sizes of $M$, and hence are bounded in terms of $\tilde{W}$. Throughout this procedure, the denominator has been contributed to only by the Plancherel density itself, which depends only on $M$. The last claim of the corollary now follows from Proposition 1. ∎

### v6

@@ 2.5 The Plancherel formula for $\mathrm{GL}_{n}$
$f(i_{M})=\sigma(i_{M})\delta^{\frac{1}{2}}_{P}(i_{M})f(1)=f(1).$
As $\delta_{P}=1$ on every compact subgroup of $P$, we have $\delta^{\frac{1}{2}}_{P}(i_{M})=1$, and $f(1)\in\sigma^{I_{M}}$. ∎
@@
Let $\pi:=\operatorname{Ind}_{P}^{G}(\nu\otimes\omega)$ be a tempered irreducible representation of $G$ induced as usual from a standard parabolic subgroup $P$ with Levi subgroup $M$. Then $\pi^{I}$ is a tempered irreducible $H$-module, and is of the form $K(u,s,\rho,q)$ for $u,s\in\mathbf{G}^{\vee}(\mathbb{C})$ and $\rho$ a representation of $\pi_{0}(Z_{G^{\vee}}(u,s))$ with $s$ compact. By Theorem 7, $K(u,s,\rho,q)$ extends to a $J$-module. By definition of Lusztig’s bijection, $\pi(f_{w})\neq 0$ only if $w$ is in the two-sided cell $\mathbf{c}=\mathbf{c}(u)$ of $W_{\mathrm{aff}}$ corresponding to $u$. On the other hand, $K(u,s,\rho,q)$ is induced from a square-integrable standard module $K_{M}(u,s,\tilde{\rho},q)$ of $H(M,I_{M})$. But now [KLDeligneLanglands, Theorem 8.3] says that $(u,s)$ must be precisely as in the statement of the proposition. Thus only such summands $(f_{w})_{M}$ in the Plancherel decomposition of $f_{w}$ are nonzero. ∎
@@ 2.8.2 Conventions on parabolic subgroups
where we view $\beta_{i}$ as restricted to $\mathfrak{a}_{P}\mathrel{\hookrightarrow}\mathfrak{a}_{G}$. Note that all the positive roots $\Sigma^{+}(\theta)$ of $(P_{\theta},N_{\theta})$ are in $N_{P}$. Denoting by $\alpha]$ the coset representing a root $\alpha$ of $G$ restricted to $P$, the positive roots in $N_{P}$ are the $\alpha_{ij}$ such that $\alpha_{ij}]=[\beta_{n_{1}+\cdots+n_{k}}]$.
@@ 2.8.3 Regularity of the trace
That is, the trace is a regular function on $(\mathbb{C}^{\times})^{k}$.
@@
When $\nu$ is non-strictly positive with respect to $M$, by Theorem 7, $\mathrm{trace}\left({\pi}\,,{f}\right)$ has poles only of the form $z_{i}^{n_{i}}=0$. The claim now follows from the $W$-invariance, and thus $\mathrm{trace}\left({\pi}\,,{f}\right)$ is a regular function on $(\mathbb{C}^{\times})^{k}$. ∎
@@
Let $d\in W_{\mathrm{aff}}$ be a distinguished involution and $\pi$ be a tempered representation induced from one of the Levi subgroups attached to the two-sided cell containing $d$ in the sense of Proposition 1. Then $\mathrm{trace}\left({\pi}\,,{f_{d}}\right)$ is constant and a natural number. In fact, the same is true for any idempotent in $J$.
@@
3 Proof of Theorem 2 for general $\mathbf{G}$, and the case of $\mathrm{GL}_{n}$
@@ 3.1 The functions $f_{w}$ for $\mathrm{GL}_{n}$
To compute with the Plancherel formula, we will need to apply the residue theorem successively in each variable $z_{i}$, and in doing so will we will need to sum over a certain tree that will track, for each variable, at which residues we evaluated. Upon integrating with respect to each variable $z_{i}$, we will have poles of the form $z_{i}=0$ or $z_{i}=(q^{ij})^{-1}z_{j}$. For example, if we have $4$ variables $z_{1},z_{2},z_{3},z_{4}$ corresponding to a Levi subgroup $\mathrm{GL}_{l_{1}}\times\mathrm{GL}_{l_{2}}\times\mathrm{GL}_{l_{3}}\times\mathrm{GL}_{l_{4}}$, then some of the summands obtained by successively applying the residue theorem are labelled by paths on the tree
@@
Given a Levi subgroup $M$ with $N+1$ blocks, a bookkeeping tree $T$ for $M$ is a rooted tree with $N$ levels such that the vertices on the $i$-th level below the root each have NiN-i child vertices, and each vertex is decorated with an equation of the form $z_{i}=0$ or $z_{i}=(q^{ij})^{-1}z_{j}$, where the index $j$ does not appear along the path from the vertex to the root, and the parent root is decorated with an equation $z_{k}=(q^{ki})^{-1}z_{i}$ for some $k$. Moreover, we require that the root be decorated with an equation of the form $z_{i}=0$ or $z_{i}=(q^{ij})^{-1}z_{j}$ for $i$ minimal. A branch of $T$ is a simple path in $T$ from the root to one of the leaves.
@@
We will show that each of the three forms of denominator that appear in the conclusion of Theorem 11 divide $P_{\mathbf{G}/\mathbf{B}}(q)$, and thus that their product divides a power of $P_{\mathbf{G}/\mathbf{B}}(q)$. The denominators of $c_{M}$ are all of the form $1+q+\cdots+q^{l_{i}-1}$, and so divide $P_{\mathbf{G}/\mathbf{B}}(q)$ as $l_{i}\leq n$ for all $i$. Note that as $l_{i_{i}}+l_{i_{k+1}}\leq n$, the leftmost denominators in (12) satisfy the conclusion of the corollary also. Finally, $Q_{rk}q^{i_{r}i_{k+1}}=q^{l_{i_{k+1}}+l_{i_{k}}+\cdots+l_{i_{r}}}$, and $R_{rk}$ is likewise always a polynomial in $q$ (as opposed to $q^{-1}$) divisible by 1q1-q. Again using that $l_{i_{k+1}}+l_{i_{k}}+\cdots+l_{i_{r}}\leq n$, we are done with the first statement.
@@
for tempered principal series. Now let $w\in\mathbf{c}_{0}$ and consider $f_{w}=\tilde{\phi}(t_{w})$. By Lemma 5 and the Satake isomorphism, there is a function $h_{w}$ in the spherical Hecke algebra such that for all principal series $\pi$,
@@
$f_{w}(1)=\int{\mathrm{trace}\left(\pi_{I}(f_{w})\right)}\,\mathrm{d}{\pi_{\mu_{I}}}=\int{\mathrm{trace}\left(\pi_{K}(h_{w})\right)}\,\mathrm{d}{\pi_{\mu_{I}}}=\frac{1}{P_{\mathbf{G}/\mathbf{B}}(q)}\int{\mathrm{trace}\left(\pi_{K}(h_{w})\right)}\,\mathrm{d}{\pi_{\mu_{K}}}=\frac{h_{w}(1)}{P_{\mathbf{G}/\mathbf{B}}(q)},$ (14)
@@
Here the rank is given by [XiI], Proposition 5.5. (In op. cit. there is the assumption of simple-connectedness, but it is easy to see that the distinguished involutions for the extended affine Weyl group $\tilde{W}(\tilde{\mathbf{G}}(F))$ of the universal cover $\tilde{\mathbf{G}}$ are distinguished involutions for $W_{\mathrm{aff}}$ using the definition in [affineII] and uniqueness of the $\{C_{w}\}$-basis, and that the lowest cell is just the lowest cell of $\tilde{W}(\tilde{\mathbf{G}}(F))$ intersected with $W_{\mathrm{aff}}$.) ∎
@@
Let $w\in W_{\mathrm{aff}}$. Then $f_{w}(1)$ is a rational function of $q$ with denominator dividing a power of the Poincaré polynomial of $\mathbf{G}$. The numerator is a Laurent polynomial $p_{1}(q^{1/2})+p_{2}(q^{-1/2})$ in $q^{1/2}$, where the degree of $p_{1}$ is bounded uniformly in terms of $W_{\mathrm{aff}}$. The denominator of $f_{w}(1)$ depends only on the two-sided cell containing $w$.
@@
where we view $\nu$ as a semisimple conjugacy class in $M_{\mathbf{c}}$, and $V(\lambda)$ is the irreducible representation of $M_{\mathbf{c}}$ of highest weight $\lambda$. Then we have that $f_{\lambda}(1)\neq 0$ only if $\lambda$ is of height 0 with respect to the basis $\varepsilon_{i}\colon\mathrm{diag}(a_{1},\dots,a_{n})\mapsto a_{i}$.
@@ 3.1.2 Proofs of Theorem 11, Lemma 7, and Corollary 3
which is one of the factors in the product $(1-Q_{r,k-1}q_{i_{r}i_{k}})(Q_{r,k-1}(q_{i_{r}i_{k}})^{-1})$. The surviving factor in the numerator at index (k1,r)(k-1,r) is then equal to $(1-Q_{r,k-1}q^{g_{i_{r}}-g_{i_{k}}})$. In short, the above factors in the denominator cancel with a numerator occurring with the same $r$-index but $k$-index one lower. Such a factor occurs whenever r<k1r<k-1 (note that this inequality does not hold when $k=1$ and $r=0$). When r=k1r=k-1, we have
@@
When evaluating (16) along a single branch, we find again that the only poles that appear are of the form $z_{i}=(q^{ij})^{-1}z_{j}$, or $z_{i}^{e_{i}}=0$ for $e_{i}<0$. We will track the effect that evaluating each successive residue has on the total degree of the integrand, and then conclude using the fact that $\int_{\mathbb{T}}z^{r}\,\mathrm{d}z=2\pi i\delta_{r,-1}$. First, observe that evaluating a residue of the form $z_{i}=(q^{ij})^{-1}z_{j}$ increases the sum of all powers by $1$; a factor $z_{j}$ is contributed to the resulting integrand. The sum of all powers is likewise increased by $1$ when evaluating the residue at a simple pole of $z_{i}^{-1}$ at 0. To compute the residue at a pole of $z_{i}^{-e_{i}}$ at 0 for $e_{i}>1$, consider the Taylor expansion at 0 of $\Gamma^{ij}$. We have
@@
$1+\frac{q^{ij}+(q^{ij})^{-1}-q_{ij}-q_{ij}^{-1}}{z_{j}}z_{i}+\frac{(q^{ij})^{2}+2+(q^{ij})^{-2}-(q_{ij}+q_{ij}^{-1})(q^{ij}+q^{-ij})}{z_{j}^{2}}z_{i}^{2}+\cdots\\ +\frac{-(q_{ij}+q_{ij}^{-1})((q^{ij})^{n-1}+\cdots+(q^{ij})^{-n+1})+(q^{ij})^{n-2}+\cdots+(q^{ij})^{-n+2}+(q^{ij})^{n}+\cdots+(q^{ij})^{-n}}{z_{j}^{n}}z_{i}^{n}\\ +\cdots.$ (19)
It is clear that the salient point of (19), that $z_{i}^{n}$ appears with a coefficient proportional to $z_{j}^{-n}$, holds also for the all the products like those in (18). Thus computing a residue at the pole of $z_{i}^{-e_{i}}$ at 0 will result in a new integrand, the total degree of which has increased by $e_{i}-e_{i}+1=1$. It now follows that after integrating with respect to N1N-1 variables, the final integral to be computed will be a constant times $(2\pi i)^{-1}\int_{\mathbb{T}}z_{N}^{e_{0}+\cdots+e_{N}+N-1}\mathrm{d}z_{N}$. This is nonzero if and only if $e_{0}+\cdots+e_{N}=-N$. Summing over all branches of a bookkeeping tree, we see that (16) is nonzero only if $e_{0}+\cdots+e_{N}=-N$. ∎
@@
The total degree of the integrand is now $e_{0}+\cdots+e_{N}+1=-(N-1)$. Therefore either $e_{i_{0}}+e_{i_{1}}+1=e_{i_{2}}=\cdots=e_{-N}=-1$, or we may again assume without loss of generality that the exponent of some $z_{i_{j}}$ is nonnegative, and proceed with evaluating (20) by integrating with respect to $z_{i_{j}}$. We may continue in this way, never having to deal with more than a simple pole at 0. (Of course, the resulting order of integration need not be the same as the order used in the proofs of Theorem 11 and Corollary 3.)
@@
for new exponents $e^{\prime}_{i_{k}}$. This and the calculations in the proof of Theorem 11 make clear that the only positive powers of $q^{1/2}$ that appear in the any denominator are those that appeared as in Theorem 11. These are controlled by the possible block sizes of $M$, and hence are bounded in terms of $W_{\mathrm{aff}}$. Throughout this procedure, the denominator has been contributed to only by the Plancherel density itself, which depends only on $M$. The last claim of the corollary now follows from Proposition 1. ∎